OpenStudy (anonymous):

find f^-1(x) if f(x) = x^3+1

OpenStudy (anonymous):

solve for y \[x^3=y-1\rightarrow x=\sqrt[3]{y-1}\rightarrow f^-1(x)=\sqrt[3]{x-1}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome :)